Initially, capacitor 50 μF is charged to voltage 80 V. This capacitor is then connected in parallel to an identical uncharged capacitor. Find the difference between the initial and final electrical energy of the system. The answer should be given in millijoules [mJ].
1
Expert's answer
2020-11-10T07:07:41-0500
"CU=C(0.5U)+C(0.5U)\\\Delta E=0.5CU^2-2(0.5)C(0.5U)^2=0.25CU^2\\\Delta E=0.25(50\cdot10^{-6})(80)^2=0.08\ J=80\ mJ"